### v1

@@ Abstract
In medicine, large language models (LLMs) commonly leverage retrieval-augmented generation (RAG) to ground outputs in up-to-date external evidence and reduce hallucinations. However, current RAG approaches focus primarily on performance improvements while overlooking evidence-based medicine (EBM) principles. This study addresses two key gaps: (1) the lack of PICO alignment between queries and retrieved evidence, and (2) the absence of evidence hierarchy considerations during reranking. We present a reusable adaptation strategy for integrating EBM into graph-based RAG, integrating the PICO framework into knowledge graph construction and retrieval, and proposing a Bayesian-inspired reranking algorithm to calibrate ranking scores by evidence grade without introducing predefined weights. We validated this framework in sports rehabilitation, a literature-rich domain currently lacking RAG systems and benchmarks. We released a knowledge graph (357,844 nodes and 371,226 edges) and a reusable benchmark of 1,637 QA pairs. The system achieved 0.830 nugget coverage, 0.819 answer faithfulness, 0.882 semantic similarity, and 0.788 PICOT match accuracy. In a 5-point Likert evaluation, five expert clinicians rated the system 4.66–4.84 across factual accuracy, faithfulness, relevance, safety, and PICO alignment. These findings demonstrate that the proposed EBM adaptation strategy substantially improves retrieval and answer quality and is readily transferable to other clinical domains; while the released resources help address the scarcity of RAG datasets in sports rehabilitation.
keywords:
Large language models, Evidence-based medicine, Knowledge graph, Retrieval-augmented generation, Sports rehabilitation
@@ 1 Introduction
Consider the following question posed to a large language model (LLM): “My child has congenital heart disease and has just undergone surgery. How should we conduct postoperative exercise rehabilitation?” As of 2024, the evidence base for rehabilitation in children with congenital heart disease remains limited, consisting primarily of observational studies[1, 2], with authoritative guidelines only recently emerging[3]. In such cases, LLMs often provide outdated or suboptimal answers due to their training data cutoff[4].
@@
In medicine, clinical evidence evolves rapidly. For example, the latest American College of Sports Medicine (ACSM) edition adopts the metabolic chronotropic reserve (MCR) to assess whether the heart rate response during exercise is appropriate[5], replacing the traditional 220 minus age formula, which recent studies have shown can carry significant errors[6, 7]. The ACSM has also updated exercise prescriptions for special populations (e.g., transgender individuals, older adults, and children), rehabilitation strategies for chronic diseases, and predictive equations[5]. Retrieval-Augmented Generation (RAG) is designed to incorporate such time-sensitive information, dynamically extending LLM knowledge to prevent obsolescence and reduce hallucinations[8]. In brief, RAG comprises three stages: the Retriever searches user-curated corpora in a vector database; the Augmenter filters, reranks, and injects results into the model’s context; and the Generator synthesizes retrieved information to produce answers with cited sources[8]. In medicine, the need for traceable evidence and timely information makes RAG a widely adopted approach for enhancing LLM reliability[9, 10].
With the rapid advancement of artificial intelligence (AI), RAG methods have been evolving rapidly. The seminal RAG framework by Lewis et al. employed a vector retriever to recall the top-$K$ text chunks from a knowledge base, which were concatenated with the user query and fed into an LLM for answer generation[8]. In 2021, Luan et al. proposed a hybrid retrieval strategy within RAG that significantly improved the reliability of downstream tasks[11]. Specifically, this strategy fuses sparse retrieval (e.g., keyword-based methods) with dense retrieval (e.g., semantic embeddings) at the recall stage, thereby enhancing retrieval performance. Subsequently, domain-specific RAG systems for medicine began to emerge. In 2023, Shi et al. introduced MKRAG, the first to incorporate a structured medical knowledge base, and conducted systematic evaluations that established a baseline for subsequent medical RAG research[12]. The following year, RAG research began to introduce domain-specific methodological innovations for medicine. For example, the $\mathrm{RAG}^{2}$ framework by Sohn et al. leverages intermediate reasoning steps generated by the LLM during answer generation to iteratively retrieve supporting evidence, thereby reducing hallucinations and enabling a traceable evidence chain[13]. Additionally, MRD-RAG by Sun et al. combines disease information trees with pseudo-medical-history indexing to develop the first multi-turn diagnostic RAG system, which substantially improves diagnostic accuracy and serves as an early prototype of Agentic RAG in medicine[14].
@@
Returning to the original user question, the patient is a child with congenital heart disease—a special population. Recently evolved RAG frameworks typically focus on performance improvements without explicitly accounting for query-specific target populations[15, 16, 17, 18], potentially generating population-mismatched answers (especially when limited corpus evidence leads to insufficient recall[19]). Crucially, most evaluation metrics cannot identify population mismatches. For example, semantic similarity uses cosine similarity to quantify differences between generated and reference answers at a macro level, making general phrasing far more influential than specific details; answer faithfulness measures only factual alignment between the answer and retrieved context, ignoring relevance to the query—retrieving passages about adults with congenital heart disease would not affect this metric[20]. Another key issue is that even if RAG incorporates the latest authoritative guidelines for the rehabilitation of children with congenital heart disease, these sources would still be weighted the same as earlier scattered observational evidence during retrieval[21], which violates the concept of a hierarchy of evidence[22]. The two issues above can be summarized as follows: Current medical RAG approaches neglect the evidence-based medicine (EBM) framework[23]. In particular, the Population–Intervention–Comparator–Outcome (PICO) framework and the hierarchy of evidence remain largely overlooked.
@@
Regarding the combination of EBM and RAG, especially the integration of the PICO framework, graph-based retrieval-augmented generation (GraphRAG) shows promising potential. This paradigm leverages knowledge graphs constructed from entity-relation-attribute triples to provide a structured hierarchy for the corpus[24]. Originally, it’s designed to enhance the multihop reasoning capabilities of LLMs for complex questions, which are particularly important in medical contexts[25]. However, this highly structured arrangement can also implicitly guide LLMs to retrieve from nodes aligned with the query, thereby potentially improving PICO alignment between questions and answers. Based on this, we propose the key question of this study:
@@
Since the pioneering work of Edge et al.[24], GraphRAG has evolved along two main trajectories: graph retrieval and graph construction[26]. The former focuses on improving retrieval efficiency and relevance, while the latter aims to enrich graph topology and enable multi-granularity retrieval. Recently, Youtu-GraphRAG, proposed by Dong et al.[26], unifies graph retrieval and graph construction through schema-bounded agentic extraction. This strategy reduces extraction noise during graph construction while enabling rapid domain adaptation through schema replacement. Essentially, it shifts the implicit retrieval constraints of traditional GraphRAG to a new paradigm of explicitly constrained retrieval by predefining the schema. In the medical domain, when we directly define the schema as PICO-related entity types, PICO mismatch between queries and retrieved chunks may be substantially reduced (Fig. 1). Accordingly, this study aims to address EBM adaptation based on the Youtu-GraphRAG framework. In particular, although the field of sports rehabilitation is resource-rich, no domain-specific RAG system or reusable benchmark currently exists. We therefore focus on this domain, curating a corpus, constructing a benchmark, and validating our framework.
@@
  • We curate high-quality corpora covering 21 common conditions in sports rehabilitation along with domain-relevant general guidelines, and construct a knowledge graph comprising 357,844 nodes and 371,226 relationships.
@@
  • We propose a Bayesian-inspired reranking algorithm that calibrates the ranking scores of a cross-encoder by incorporating both semantic and evidence-grade information of candidate chunks without introducing predefined weights, enabling explicit consideration of evidence hierarchy during reranking.
@@
  • We present Sports Rehabilitation RAG (SR-RAG), which integrates the EBM framework into GraphRAG pipeline through a generalizable adaptation strategy, with sports rehabilitation as the validation domain.

### v2

@@ Abstract
Current medical retrieval-augmented generation (RAG) approaches overlook evidence-based medicine (EBM) principles, leading to two key gaps: (1) the lack of PICO alignment between queries and retrieved evidence, and (2) the absence of evidence hierarchy considerations during reranking. We present SR-RAG, an EBM-adapted GraphRAG framework that integrates the PICO framework into knowledge graph construction and retrieval, and proposes Bayesian Evidence Tier Reranking (BETR) to calibrate ranking scores by evidence grade without predefined weights. Validated in sports rehabilitation, we release a knowledge graph (357,844 nodes, 371,226 edges) and a benchmark of 1,637 QA pairs. SR-RAG achieves 0.812 evidence recall@10, 0.830 nugget coverage, 0.819 answer faithfulness, 0.882 semantic similarity, and 0.788 PICOT match accuracy, substantially outperforming five baselines. Five expert clinicians rated the system 4.66–4.84 on a 5-point Likert scale, and system rankings are preserved on a human-verified gold subset ($n=80$).
From Evidence-Based Medicine to Knowledge Graph: Retrieval-Augmented Generation for Sports Rehabilitation and a Domain Benchmark
Jinning Zhang1,2∗   Jie Song1,2∗   Wenhui Tu1,2‡   Zecheng Li1,2‡ Jingxuan Li3   Jin Li4   Xuan Liu5   Taole Sha7   Zichen Wei6   Yan Li1,2† 1Beijing Key Laboratory of Sports Performance and Skill Assessment, Beijing Sport University, Beijing 100084, China 2Department of Exercise Physiology, School of Sport Science, Beijing Sport University, Beijing 100084, China 3School of Sport Medicine and Physical Therapy, Beijing Sport University, Beijing 100084, China 4Rehabilitation Center, Beijing Rehabilitation Hospital, Beijing 100144, China 5Optum Care Washington, Everett, Washington 98201, USA 6School of Management, Beijing Sport University, Beijing 100084, China 7Department of Statistics and Actuarial Science, The University of Hong Kong, Hong Kong, China Equal contribution  Equal contribution as second authors  Corresponding author: bsuliyan@bsu.edu.cn
@@ 1 Introduction
Consider the following question posed to a large language model (LLM): “My child has congenital heart disease and has just undergone surgery. How should we conduct postoperative exercise rehabilitation?” As of 2024, the evidence base for rehabilitation in children with congenital heart disease remains limited, consisting primarily of observational studies (Barbazi et al., 2025; Ubeda Tikkanen et al., 2023), with authoritative guidelines only recently emerging (Interamerican Society of Cardiology (2024), SIAC). In such cases, LLMs often provide outdated or suboptimal answers due to their training data cutoff (Hager et al., 2024).
@@
In medicine, clinical evidence evolves rapidly. For example, the latest American College of Sports Medicine (ACSM) edition adopts the metabolic chronotropic reserve (MCR) to assess whether the heart rate response during exercise is appropriate (American College of Sports Medicine, 2025), replacing the traditional 220 minus age formula, which recent studies have shown can carry significant errors (Almaadawy et al., 2024; Lauer et al., 1999). Retrieval-Augmented Generation (RAG) is designed to incorporate such time-sensitive information, dynamically extending LLM knowledge to prevent obsolescence and reduce hallucinations (Lewis et al., 2020). In medicine, the need for traceable evidence and timely information makes RAG a widely adopted approach for enhancing LLM reliability (Yang et al., 2025; Béchard and Ayala, 2024).
@@
Returning to the original user question, the patient is a child with congenital heart disease—a special population. Recently evolved RAG frameworks typically focus on performance improvements without explicitly accounting for query-specific target populations (Gupta et al., 2024; Xiong et al., 2024; Asai et al., 2024; Yan et al., 2024), potentially generating population-mismatched answers. Crucially, most evaluation metrics cannot identify population mismatches. For example, semantic similarity uses cosine similarity to quantify differences between generated and reference answers at a macro level, making general phrasing far more influential than specific details; answer faithfulness measures only factual alignment between the answer and retrieved context, ignoring relevance to the query (Es et al., 2024). Another key issue is that even if RAG incorporates the latest authoritative guidelines for the rehabilitation of children with congenital heart disease, these sources would still be weighted the same as earlier scattered observational evidence during retrieval (Sun et al., 2025b), which violates the concept of a hierarchy of evidence (Guyatt et al., 2008). The two issues above can be summarized as follows: Current medical RAG approaches neglect the evidence-based medicine (EBM) framework (Sackett et al., 1996). In particular, the Population–Intervention–Comparator–Outcome (PICO) framework and the hierarchy of evidence remain largely overlooked.
@@
Regarding the combination of EBM and RAG, especially the integration of the PICO framework, graph-based retrieval-augmented generation (GraphRAG) shows promising potential. This paradigm leverages knowledge graphs constructed from entity-relation-attribute triples to provide a structured hierarchy for the corpus (Edge et al., 2024). Originally, it is designed to enhance the multihop reasoning capabilities of LLMs for complex questions, which are particularly important in medical contexts (Cabello et al., 2024). However, this highly structured arrangement can also implicitly guide LLMs to retrieve from nodes aligned with the query, thereby potentially improving PICO alignment between questions and answers. Based on this, we propose the key question of this study:
@@
Since the pioneering work of Edge et al. (2024), GraphRAG has evolved along two main trajectories: graph retrieval and graph construction (Dong et al., 2026). The former focuses on improving retrieval efficiency and relevance, while the latter aims to enrich graph topology and enable multi-granularity retrieval. Recently, Youtu-GraphRAG, proposed by Dong et al. (2026), unifies graph retrieval and graph construction through schema-bounded agentic extraction. This strategy reduces extraction noise during graph construction while enabling rapid domain adaptation through schema replacement. Essentially, it shifts the implicit retrieval constraints of traditional GraphRAG to a new paradigm of explicitly constrained retrieval by predefining the schema. In the medical domain, when we directly define the schema as PICO-related entity types, PICO mismatch between queries and retrieved chunks may be substantially reduced (Fig. 1). Accordingly, this study aims to address EBM adaptation based on the Youtu-GraphRAG framework. In particular, although the field of sports rehabilitation is resource-rich, no domain-specific RAG system or reusable benchmark currently exists. We therefore focus on this domain, curating a corpus, constructing a benchmark, and validating our framework.
@@
  • We curate a sports rehabilitation corpus (21 conditions) and construct a knowledge graph comprising 357,844 nodes and 371,226 edges.
@@
  • We propose BETR, a Bayesian-inspired reranking algorithm that calibrates ranking scores by learning evidence-grade biases from data without predefined weights.
@@
  • We present SR-RAG, integrating the EBM framework into GraphRAG through a generalizable adaptation strategy.